OpenStudy (the-simple-diamond):

Jane bennett can invest $20000 at either 5.5% coumpounded monthly for 1 year (amount of $1 investment is 1.156408) Or 6.25% compounded quarterly for 1 uear ( amount of $1 investment is 1.06398) What is the effective annual yield of the better onvestment?

OpenStudy (wolf1728):

Compound Interest= (1 + rate/n)^n where n is the number of times it is compounded for 5.5% (1 + .055/12)^12 = (1.0045833333)^12 = 1.0564078604 per year for 6.25% compounded quarterly (1 + .0625/4)^4 = (1.015625)^4 = 1.0639801621 per year Which is better?

OpenStudy (the-simple-diamond):

1.0639801621

OpenStudy (the-simple-diamond):

Nvm I figured it out It was 6.398%